Devices that have been offline should now have network service. Any device still unable to access network resources should be rebooted. If you are still having difficulty, please report it to the CTS Helpdesk: 315-267-4444 If no one is available to take your call, leave a message with contact information and you will be contacted.
CTS continues to investigate why some computers/devices are not receiving network services.
CTS has reports from several areas that the network/internet is not accessible. We are investigating.
The network outage that occurred just before Noon has been resolved. If you are still experiencing difficulty getting online, please restart your device. If that does not resolve the problem, contact the CTS Helpdesk at 315-267-4444. If no one is available to answer your call, leave a message with contact information and someone will return your call.
SUNY will be performing maintenance on the SPSS License servers on Wednesday, August 2, 2017 from 4:00 A. M. to 7:00 A. M. SPSS on campus-managed computers may be unavailable during those hours.
ITEC will be performing OS and Firmware updates on Wednesday July 26th from midnight until 7:00am. During this time Banner/BearPAWS/BearDeN and SPSS will be unavailable.
Oracle Business Intelligence is back online and OK to use.
Oracle Business Intelligence is experiencing a critical error and is unavailable for account detail viewing. The problem has been reported to SUNY. We will post again when the problem is resolved.
Helios is now up.
If you don’t see the shares you expect, please try rebooting your computer. If that doesn’t work, please report the situation to the Helpdesk at extension 4444 or email@example.com.
Thank you for your patience while work was in progress.
Helios will be rebooted in one minute. Save your work on helios.
Users attempting to access helios are finding that not all of their folders are accessible. The underlying problem has been resolved, but to restore helios to full functionality it is necessary to reboot helios.
This will occur at 10:15am (in about 10 minutes). Please save any documents that came from or are going to be saved to helios, to prevent risk of corruption of those documents when the reboot occurs.
At approximately 8:30 this evening a core electrical supply failed in one of our network operations centers resulting in phone and data network interruption to most of the south end of campus. CTS personnel responded and were able to re-route power to affected systems, restoring service shortly after 10:00. If your phone does not work and/or your computer is not online, find a way to report it to the CTS Helpdesk: 315-267-4444. If no one is available to take your call, leave a message with contact information and someone will follow up with you.
ITEC will be performing OS and Firmware updates on Wednesday July 19th from midnight until 7:00am. During this time Banner/BearPAWS/BearDeN will be unavailable.
SUNY’s Information Technology Exchange Center (ITEC), which hosts the license managers for SPSS, will be performing updates on their servers this Wednesday, July 19th, 2017 between Midnight and 7:00am. Users of SPSS may find it unavailable during that time.
ITEC will be performing OS and Firmware updates on Wednesday July 26th from midnight until 7:00am. During this time Banner/BearPAWS/BearDeN will be unavailable.